Calculate the minimum velocity with which water of coefficient of viscosity 0.001Nsm^-2 flows through a tube of radius 6×10^-4 m, without turbulence being produced.Reynolds number is 1000. Find the rate of flow of water through the tube at this velocity and the pressure is required to maintain it if the length of the tube is 0.2m

1
Expert's answer
2021-11-19T10:48:17-0500

"Re=\frac{2\rho vr}{\eta},\implies"

"v=\frac{\eta Re}{2\rho r}=0.83~\frac ms,"

"Q=vS=\pi r^2 v=942~\frac{mm^3}s."